The following table represents three firms, the amount they pollute, and the cost to reduce their pollution by one ton. Use the information to answer the following questions.
Firm Pollution(tons) Cost/ton
A 50 $20
B 100 $25
C 60 $10
5. Environmental Economists have estimated that the firms are polluting too much. To
achieve the socially optimal level of pollution we need to reduce pollution by 60 tons. If
we use a command and control policy that forces each firm to reduce by the same
amount, what will be the cost of 60 tons of pollution reduction?
6. Being a student of Professor Hayes’ Econ course, you suggest that using tradeable
permits can achieve the same goal at a lower cost. How many permits should be issued to
each firm to reduce pollution by 60 tons?
7. Using tradeable permits, assume that B buys permits from C for $10 each. What is the
cost of pollution reduction after they permits are traded?
1
Expert's answer
2019-03-14T18:25:50-0400

Firm Pollution(tons) Cost/ton

A 50 $20

B 100 $25

C 60 $10

5. If we use a command and control policy that forces each firm to reduce by the same

amount, then the cost of 60 tons of pollution reduction is: 20*20 + 20*25 + 20*10 = $1100.

6. As 210 tons are produced in total, then 210 - 60 = 150 permits should be issued to

each firm to reduce pollution by 60 tons.

7. Using tradeable permits, assume that B buys permits from C for $10 each. The

cost of pollution reduction after the permits are traded is: $10*60 = $600.